# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ A.2. Computational Scaling of the A2DMRG Algorithm 2
As emphasized in Section 3.1, each global iteration of the A2DMRG Algorithm 2 essentially consists of four steps, the first of which is a sequence of LQ decompositions to designed to obtain a $j$-orthogonal copy of the initial tensor train decomposition $\mathbb{Y}_{d}^{(n)}$ for each $j\in\{d,d-1,\ldots,2\}$. The total computational cost of this step is therefore
[…]
The next step in the A2DMRG Algorithm 2 consists of performing $d$ one-site DMRG micro-iterations. As discussed in detail in Section A.1 above, each one-site DMRG micro-iteration in the present setting requires the construction of the left and right environments. Note that this is somewhat different from the classical one-site DMRG algorithm in which, during a left-to-right sweep for instance, only the right environment needs to be constructed from scratch at each micro-iteration with the left environment being obtained by a cheaper update step. Regardless, assuming that a maximum of $K\in\mathbb{N}$ matrix-vector products are required for each one-site DMRG micro-iterations, the total cost of this step of the additive one-site DMRG algorithm is (recall Estimate (32))
[…]
Turning now to the third step, i.e., the second-level minimization in the A2DMRG Algorithm 2, we see that we are required to compute the lowest eigenvalue of the symmetric eigenvalue problem (18). To do so, we must first compute the singular value decomposition of the mass matrix $\widehat{\mathbb{S}}\in\mathbb{R}^{d\times d}$ defined as
$\forall i,j,\in\{1,\ldots,d\}\colon\qquad[\widehat{\mathbb{S}}]_{i,j}=\big{\langle}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})\big{\rangle},$
where $\widetilde{\mathbb{Y}}_{\ell}^{(n+1)},\ell\in\{1,\ldots,d\}$ is the tensor train decomposition obtained by performing a one-site DMRG micro-iteration to update the $\ell^{\rm th}$ component of the initialization $\widetilde{\mathbb{Y}}_{\ell}^{(n)}$. It follows that each entry of the mass matrix $\widehat{\mathbb{S}}$ requires computing an inner product between two tensors possessing TT decompositions and therefore carries a computational cost of $\mathcal{O}(dnr^{3})$. Consequently, the total cost of constructing the symmetric matrix $\widehat{\mathbb{S}}$ and computing its singular value decomposition scales as
@@
Note that although this cost, a priori, scales cubically in the tensor order $d$, it does not involve contractions with the underlying tensor operator $\mathbb{A}$. Consequently, if the separation ranks of $\mathbb{A}$ scale (at least) quadratically with $d$ (which is the case for the second-quantized Hamiltonian in quantum chemistry [17, Section 3.2]), then the total cost of this step is comparable with the cost of constructing a left or right environment in the course of the DMRG micro-iterations described above.
Next, let us consider the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d\times d}$ appearing in the eigenvalue problem (18), and defined as
$\forall i,j,\in\{1,\ldots,d\}\colon\qquad[\widehat{\mathbb{A}}]_{i,j}=\big{\langle}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)}),\mathbb{A}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})\big{\rangle}.$
@@
  • We can appeal to a direct computation of $\widehat{\mathbb{A}}$. Following the same arguments presented in Section A.1, we see that computing each entry of $\widehat{\mathbb{A}}$ carries a computational cost of $\mathcal{O}(dn^{2}r^{3}R)$ leading to a total matrix construction cost of
@@
[…]
@@
    $\mathbb{R}^{p}\ni\mathbb{z}=(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}^{+}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}\mathbb{y}\qquad\text{for arbitrary vectors }\mathbb{y}\in\mathbb{R}^{p}.$
    Here, $p\in\mathbb{N}$ denotes the numerical rank of the mass matrix $\widehat{\mathbb{S}}$, and the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}\in\mathbb{R}^{p\times p}$ and left-orthogonal matrix $\widehat{\mathbb{U}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d\times p}$ are obtained from the singular value decomposition of $\widehat{\mathbb{S}}$ as described in Section 3.1.
@@
    $\mathbb{z}_{i}$ $=\sum_{j=1}^{d}\sum_{k=1}^{d}\sum_{\ell=1}^{p}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}]_{i,j}[\widehat{\mathbb{A}}]_{j,k}[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell}^{-1/2}\mathbb{y}_{\ell}$
@@
    $=\sum_{j=1}^{d}\sum_{k=1}^{d}\sum_{\ell=1}^{p}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}]_{i,j}\left\langle\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\mathbb{A}\tau(\widetilde{\mathbb{Y}}_{k}^{(n+1)})\right\rangle[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell}^{-1/2}\mathbb{y}_{\ell}$
@@
    $=\left\langle\sum_{j=1}^{d}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}]_{i,j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\mathbb{A}\sum_{k=1}^{d}\sum_{\ell=1}^{p}[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell}^{-1/2}\mathbb{y}_{\ell}\tau(\widetilde{\mathbb{Y}}_{k}^{(n+1)})\right\rangle$
    (35) $=\left\langle\sum_{j=1}^{d}{\mathbb{t}}_{j}\tau(\widetilde{\mathbb{Y}})_{j}^{(n+1)},\mathbb{A}\sum_{k=1}^{d}\mathbb{w}_{k}\mathbb{y}_{\ell}\tau(\widetilde{\mathbb{Y}}_{k}^{(n+1)})\right\rangle,$
    where we have introduced the vectors $\mathbb{t},\mathbb{w}\in\mathbb{R}^{d}$ given by
    $\forall j\in\{1,\ldots,d\}\colon\hskip 2.84526pt\mathbb{t}_{j}=[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}]_{i,j}~{}\text{ and }~{}\mathbb{w}_{j}=\sum_{\ell=1}^{p}[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell}^{-1/2}\mathbb{y}_{\ell}.$
@@
    In other words, the $i^{\rm th}$ entry of the output vector $\mathbb{z}$ can be expressed in terms of an inner product involving $\mathbb{A}$ and two different linear combinations of the tensor train decompositions $\{\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})\}_{j=1}^{d}$. It is well known that a linear combination of $d$ tensor train decompositions with TT ranks bounded by $r$ can itself be expressed as a tensor train decomposition with TT ranks bounded by $dr$. We claim that in fact, for any vectors $\mathbb{t},\mathbb{w}\in\mathbb{R}^{d}$, each of the linear combinations
    $\sum_{j=1}^{d}{\mathbb{t}}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\quad\sum_{k=1}^{d}\mathbb{w}_{k}\mathbb{y}_{\ell}\tau(\widetilde{\mathbb{Y}}_{k}^{(n+1)}),$
@@
    possesses a tensor train decomposition with TT ranks bounded by $2r$. To see this, recall that, by construction, each $\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$ is of the form
    $\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})=\big{(}\mathbb{U}_{1}^{(n)},\ldots\mathbb{U}_{j-1}^{(n)},\delta\mathbb{U}_{j}^{(n+1)},\mathbb{V}_{j+1}^{(n)},\ldots\mathbb{V}_{d}^{(n)}\big{)}$
@@
    where the component $\delta\mathbb{U}_{j}^{(n+1)}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ has been obtained through the $j^{\rm th}$ one-site DMRG micro-iteration applied to the initial $j$-orthogonal tensor train iterate $\tau({\mathbb{Y}})_{j}^{(n)}$. It therefore follows from a direct calculation that for all $x_{j}\in\{1,\ldots,n_{j}\},~{}j=1,\ldots,d$ we can write
    (36) $\Big{(}\sum_{j=1}^{d}\mathbb{t}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})\Big{)}(x_{1},\ldots,x_{d})$ $=\mathbb{W}_{1}(x_{1})\mathbb{W}_{2}(x_{2})\ldots\mathbb{W}_{d}(x_{d}),$
@@
    (37) $\begin{split}\mathbb{W}_{1}(x_{1})&=\begin{bmatrix}\mathbb{t}_{1}\delta\mathbb{U}^{(n+1)}_{1}(x_{1})&\mathbb{U}_{1}^{(n)}(x_{1})\end{bmatrix}\in\mathbb{R}^{r_{0}\times 2r_{1}},\\[5.0pt] \mathbb{W}_{j}(x_{j})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{j}(x_{j})&0\\ \mathbb{t}_{j}\delta\mathbb{U}^{(n+1)}_{j}(x_{j})&\mathbb{U}^{(n+1)}_{j}(x_{j})\\ \end{bmatrix}\in\mathbb{R}^{2r_{j-1}\times 2r_{j}}\quad\forall j\in\{2,\ldots,d-1\},\\ \mathbb{W}_{d}(x_{d})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{d}(x_{d})\\ \mathbb{t}_{d}\delta\mathbb{U}_{d}^{(n+1)}(x_{d})\end{bmatrix}\in\mathbb{R}^{2r_{d-1}\times r_{d}}.\end{split}$
    Consequently, the tensor $\sum_{j=1}^{d}\mathbb{t}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$ has tensor train decomposition $\big{(}\mathbb{W}_{1},\ldots\mathbb{W}_{d}\big{)}$ with components defined through Equation (37). A similar calculation of course also works when the vector $\mathbb{t}$ is replaced with $\mathbb{w}$.
    Returning now to Equation (35), we see that the $i^{\rm th}$ entry $\mathbb{z}_{i}$ of our sought-after matrix-vector product can be obtained by performing a tensor contraction of the type described in Section A.1 above. Assuming therefore, that $K^{\prime}\in\mathbb{N}$ Krylov iterations are required to solve the eigenvalue problem (18), we arrive at the following total computational cost of this third step of the additive one-site DMRG algorithm 2:
[…]
@@
    In other words, the use of a Krylov solver allows the reduction of the $d^{3}$ scaling cost in the construction of $\widehat{\mathbb{A}}$ to a $d^{2}$ scaling cost. Note that this cost can be further reduced to linear scaling in $d$ per processor provided that $d$ independent processors are available.
It now remains to consider the computational cost of the fourth and final step in the additive one-site DMRG algorithm, namely the TT rounding step. To do so, we simply appeal once again to Equations (36)-(37) to deduce that any linear combination $\sum_{j=1}^{d}{\mathbb{d}}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$ of tensor possesses a tensor train decompositions with TT ranks bounded by $2r$. The TT rounding algorithm of Oseledets [33] therefore requires a total computational cost
[…]
and we thus conclude that each global iteration of the additive algorithm 2– assuming the use of a Krylov solver in the third step– has computational cost
[…]
@@
Before proceeding to a study of the two-site A2DMRG Algorithm 2, let us conclude this discussion by emphasizing that although the computational costs of the one-site DMRG and one-site A2DMRG algorithms are ostensibly similar, A2DMRG is well-suited for parallelization. Indeed, assuming that we have at hand $d$ processors (or $d(d+1)/2$ processors if the matrices in the second-level minimization are constructed explicitly), we can distribute the workload of the computationally intensive portion of the additive algorithm 2, namely, steps two and three, to these processors and arrive at the following cost per processor for a single iteration:
(39) ${\rm Cost}_{\rm 1site,compute}^{\rm add,pp}=\mathcal{O}\big{(}nr^{3}R(K+d)\big{)}+\mathcal{O}\big{(}dnr^{3}RK^{\prime}\big{)}+\mathcal{O}(d^{2}nr^{3}).$
In the specific case of quantum chemistry applications, i.e., when the energy functional functional $\mathcal{J}$ to be minimized is of the form (2)
$\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{\langle\mathbb{U},\mathbb{H}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}},$
with $\mathbb{H}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ denoting a second-quantized molecular Hamiltonian [17, Section 3.2], it is well-known that $R=\mathcal{O}(d^{2})$. Consequently, the cost per processor for a single iteration scales in this case as
(40) ${\rm Cost}_{\rm 1site,compute}^{\rm add,pp}=\mathcal{O}\big{(}n^{2}r^{3}R(K+d)\big{)}+\mathcal{O}\big{(}dn^{2}r^{3}RK^{\prime}\big{)}.$

### v2

@@ A.2. Computational Scaling of the A2DMRG Algorithm 2
As emphasized in Section 3.1, each global iteration of the A2DMRG Algorithm 2 essentially consists of four steps, the first of which is a sequence of LQ decompositions designed to obtain a $j$-orthogonal copy of the initial tensor train decomposition $\mathbb{U}^{(n),d}$ for each $j\in\{d,d-1,\ldots,1\}$. The total computational cost of this step is therefore
${\rm Cost}^{\rm 1site}_{\rm Step\;1}=\mathcal{O}(dnr^{3}).$
The next step in the A2DMRG Algorithm 2 consists of performing $d$ one-site DMRG micro-iterations. As discussed in detail in Section A.1 above, each one-site DMRG micro-iteration in the present setting requires the construction of the left and right environments. Note that this is somewhat different from the classical one-site DMRG algorithm in which, during a left-to-right half-sweep for instance, only the right environment needs to be constructed from scratch at each micro-iteration with the left environment being obtained by a cheaper update step. Regardless, assuming that a maximum of $K\in\mathbb{N}$ matrix-vector products are required for each one-site DMRG micro-iterations, the total cost of this step of the one-site A2DMRG algorithm is (recall Estimate (32))
${\rm Cost}^{\rm 1site}_{\rm Step\;2}=\mathcal{O}\big(dnr^{3}R(K+d)\big).$
Turning now to the third step, i.e., the second-level minimization in the A2DMRG Algorithm 2, we see that we are required to compute the lowest eigenvalue of the symmetric eigenvalue problem (18). To do so, we must first compute the singular value decomposition of the mass matrix $\widehat{\mathbb{S}}\in\mathbb{R}^{d+1\times d+1}$ defined as
$\forall i,j,\in\{1,\ldots,d+1\}\colon\qquad[\widehat{\mathbb{S}}]_{i,j}=\big\langle\tau(\widetilde{\mathbb{U}}^{(n+1),{i-1}}),\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\big\rangle,$
where $\widetilde{\mathbb{U}}^{(n+1),\ell},\ell\in\{1,\ldots,d\}$ is the tensor train decomposition obtained by performing a one-site DMRG micro-iteration to update the $\ell^{\rm th}$ component of the previous global iterate $\widetilde{\mathbb{U}}^{(n+1),0}$. It follows that each entry of the mass matrix $\widehat{\mathbb{S}}$ requires computing an inner product between two tensors possessing TT decompositions and therefore carries a computational cost of $\mathcal{O}(dnr^{3})$. Consequently, the total cost of constructing the symmetric matrix $\widehat{\mathbb{S}}$ and computing its singular value decomposition scales as
@@
Note that although this cost, a priori, scales cubically in the tensor order $d$, it does not involve contractions with the underlying tensor operator $\mathbb{A}$. Consequently, if the separation ranks of $\mathbb{A}$ scale (at least) quadratically with $d$ (which is the case for the second-quantized Hamiltonian in quantum chemistry [19, Section 3.2]), then the total cost of this step is comparable with the cost of constructing a left or right environment in the course of the DMRG micro-iterations described above.
Next, let us consider the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d+1\times d+1}$ appearing in the eigenvalue problem (18), and defined as
$\forall i,j,\in\{1,\ldots,d+1\}\colon\qquad[\widehat{\mathbb{A}}]_{i,j}=\big\langle\tau(\widetilde{\mathbb{U}}^{(n+1),{i-1}}),\mathbb{A}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\big\rangle.$
@@
  • We can appeal to a direct computation of $\widehat{\mathbb{A}}$. Following the same arguments presented in Section A.1, computing each entry of $\widehat{\mathbb{A}}$ carries a computational cost of $\mathcal{O}(dnr^{3}R)$ leading to a total matrix construction cost of
@@
    ${\rm Cost}^{\rm 1site}_{\rm Step\;3,direct}=\mathcal{O}\Big(\frac{d(d+1)}{2}dnr^{3}\Big)+\mathcal{O}\Big(\frac{d(d+1)}{2}dnRr^{3}\Big)+\mathcal{O}(d^{3}).$
@@
    $\mathbb{R}^{p}\ni\mathbb{z}=(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}^{+}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}\mathbb{y}\qquad\text{for arbitrary vectors }\mathbb{y}\in\mathbb{R}^{p}.$
    Here, $p\in\mathbb{N}$ denotes the numerical rank of the mass matrix $\widehat{\mathbb{S}}$, and the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}\in\mathbb{R}^{p\times p}$ and left-orthogonal matrix $\widehat{\mathbb{V}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d+1\times p}$ are obtained from the singular value decomposition of $\widehat{\mathbb{S}}$ (see Section 3.1).
@@
    $\mathbb{z}_{i}=$ $\sum_{j=1}^{d+1}\sum_{k=1}^{d+1}\sum_{\ell=1}^{p}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}]_{i,j}[\widehat{\mathbb{A}}]_{j,k}[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell,\ell}^{-1/2}\mathbb{y}_{\ell}$
@@
    $=$ $\sum_{j=1}^{d+1}\sum_{k=1}^{d+1}\sum_{\ell=1}^{p}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}]_{i,j}\left\langle\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}}),\mathbb{A}\tau(\widetilde{\mathbb{U}}^{(n+1),{k-1}})\right\rangle[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell,\ell}^{-1/2}\mathbb{y}_{\ell}$
@@
    $=$ $\left\langle\sum_{j=1}^{d+1}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}]_{i,j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}}),\mathbb{A}\sum_{k=1}^{d+1}\sum_{\ell=1}^{p}[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell,\ell}^{-1/2}\mathbb{y}_{\ell}\tau(\widetilde{\mathbb{U}}^{(n+1),{k-1}})\right\rangle$
    (35) $=$ $\left\langle\sum_{j=1}^{d+1}{\mathbb{t}}^{(i)}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}}),\mathbb{A}\sum_{k=1}^{d+1}\mathbb{w}_{k}\tau(\widetilde{\mathbb{U}}^{(n+1),{k-1}})\right\rangle,$
    where we have introduced the vectors $\mathbb{t}^{(i)},\mathbb{w}\in\mathbb{R}^{d+1}$ given by
    $\forall j\in\{1,\ldots,d+1\}\colon\hskip 2.84526pt\mathbb{t}^{(i)}_{j}=[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{i,i}^{-1/2}[(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}]_{i,j},\hskip 2.84526pt\mathbb{w}_{j}=\sum_{\ell=1}^{p}[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{k,\ell}[\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}]_{\ell,\ell}^{-1/2}\mathbb{y}_{\ell}.$
@@
    In other words, the $i^{\rm th}$ entry of the output vector $\mathbb{z}$ can be expressed in terms of an inner product involving $\mathbb{A}$ and two different linear combinations of the tensor train decompositions $\{\tau(\widetilde{\mathbb{U}}^{(n+1),j})\}_{j=0}^{d}$. It is well known that a linear combination of $d$ tensor train decompositions with TT ranks bounded by $r$ can itself be expressed as a tensor train decomposition with TT ranks bounded by $dr$. We claim that in fact, for any vectors $\mathbb{t}^{(i)},\mathbb{w}\in\mathbb{R}^{d}$, each of the linear combinations
    $\sum_{j=1}^{d+1}{\mathbb{t}}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}}),\quad\sum_{k=1}^{d+1}\mathbb{w}_{k}\tau(\widetilde{\mathbb{U}}^{(n+1),{k-1}}),$
@@
    possesses a tensor train decomposition with TT ranks bounded by $2r$. To see this, recall that, by construction, each $\widetilde{\mathbb{U}}^{(n+1),j},~j\in\{1,\ldots,d\}$ is of the form
    $\widetilde{\mathbb{U}}^{(n+1),j}=\big(\mathbb{U}_{1}^{(n)},\ldots\mathbb{U}_{j-1}^{(n)},\delta\mathbb{U}_{j}^{(n+1)},\mathbb{V}_{j+1}^{(n)},\ldots\mathbb{V}_{d}^{(n)}\big),$
@@
    where the component $\delta\mathbb{U}_{j}^{(n+1)}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ has been obtained through the $j^{\rm th}$ one-site DMRG micro-iteration applied to the initial $j$-orthogonal tensor train iterate $\tau({\mathbb{U}}^{(n),j})$. Since $\tau({\mathbb{U}}^{(n+1),0})=\tau({\mathbb{U}}^{(n),j})$ for all $j\in\{1,\ldots,d\}$, we can, e.g., also take the right-orthogonal tensor train decomposition of $\tau({\mathbb{U}}^{(n+1),0})$, i.e.,
    $\tau({\mathbb{U}}^{(n+1),0})=\tau\big({\delta W_{1}}^{(n)},\mathbb{V}_{2}^{(n)}\ldots\mathbb{V}_{d}^{(n)}\big).$
    It follows from a direct calculation that for all $x_{\ell}\in\{1,\ldots,n_{\ell}\},~\ell=1,\ldots,d$ we can write
    (36) $\Big(\sum_{j=1}^{d+1}\mathbb{t}^{(i)}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\Big)(x_{1},\ldots,x_{d})$ $=\mathbb{W}_{1}(x_{1})\mathbb{W}_{2}(x_{2})\ldots\mathbb{W}_{d}(x_{d}),$
@@
    (37) $\begin{split}\mathbb{W}_{1}(x_{1})&=\begin{bmatrix}\mathbb{t}_{1}^{(i)}{\delta W_{1}}^{(n)}(x_{1})+\mathbb{t}_{2}^{(i)}\delta\mathbb{U}^{(n+1)}_{1}(x_{1})&\mathbb{U}_{1}^{(n)}(x_{1})\end{bmatrix}\in\mathbb{R}^{r_{0}\times 2r_{1}},\\[5.0pt] \mathbb{W}_{\ell}(x_{\ell})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{\ell}(x_{\ell})&0\\ \mathbb{t}_{\ell+1}^{(i)}\delta\mathbb{U}^{(n+1)}_{\ell}(x_{\ell})&\mathbb{U}^{(n+1)}_{\ell}(x_{\ell})\\ \end{bmatrix}\in\mathbb{R}^{2r_{\ell-1}\times 2r_{\ell}}\hskip 2.84526pt\forall\ell\in\{2,\ldots,d-1\},\\ \mathbb{W}_{d}(x_{d})&=\begin{bmatrix}\mathbb{V}^{(n+1)}_{d}(x_{d})\\ \mathbb{t}_{d+1}^{(i)}\delta\mathbb{U}_{d}^{(n+1)}(x_{d})\end{bmatrix}\in\mathbb{R}^{2r_{d-1}\times r_{d}}.\end{split}$
    Thus, the tensor $\sum_{j=1}^{d+1}\mathbb{t}^{(i)}_{j}\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})$ has a tensor train decomposition $\big(\mathbb{W}_{1},\ldots\mathbb{W}_{d}\big)$ with components defined through Equation (37). A similar calculation of course also works when the vector $\mathbb{t}^{(i)}$ is replaced with $\mathbb{w}$.
    Returning now to Equation (35), we see that the $i^{\rm th}$ entry $\mathbb{z}_{i}$ of our sought-after matrix-vector product can be obtained by performing a tensor contraction of the type described in Section A.1 above. Assuming therefore, that $K^{\prime}\in\mathbb{N}$ Krylov iterations are required to solve the eigenvalue problem (18), we arrive at the following total computational cost of this third step of the one-site A2DMRG algorithm 2:
    ${\rm Cost}_{\rm Step\;3,Krylov}^{\rm 1site}=$ $\mathcal{O}\big(d^{2}nr^{3}RK^{\prime}\big)+{\rm Cost}_{\mathbb{S},\rm SVD}$
@@
    In other words, the use of a Krylov solver allows the reduction of the $d^{3}$ scaling cost in the construction of $\widehat{\mathbb{A}}$ to a $d^{2}$ scaling cost. This cost can be further reduced to linear scaling in $d$ per processor provided that $d$ independent processors are available.
It now remains to consider the computational cost of the fourth and final step in the one-site A2DMRG algorithm, namely the TT rounding step. To do so, we simply appeal once again to Equations (36)-(37) to deduce that any linear combination $\sum_{j=1}^{d}{\mathbb{d}}_{j}\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)})$ of tensor possesses a tensor train decompositions with TT ranks bounded by $2r$. The TT rounding algorithm of Oseledets [35] therefore requires a total computational cost
${\rm Cost}_{\rm Step\;4}^{\rm 1site}=\mathcal{O}(dnr^{3}),$
and we thus conclude that each global iteration of the A2DMRG Algorithm 2– assuming the use of a Krylov solver in the third step– has computational cost
${\rm Cost}_{\rm 1site,compute}^{\rm A2DRMG}$ $={\rm Cost}^{\rm 1site}_{\rm Step\;1}+{\rm Cost}^{\rm 1site}_{\rm Step\;2}+{\rm Cost}^{\rm 1site}_{\rm Step\;3,Krylov}+{\rm Cost}^{\rm 1site}_{\rm Step\;4}$
@@
Before proceeding to a study of the two-site A2DMRG Algorithm 2, let us conclude this discussion by emphasizing that although the computational costs of the one-site DMRG and one-site A2DMRG algorithms are ostensibly similar, A2DMRG is well-suited for parallelization. Indeed, assuming that we have at hand $d$ processors (or $d(d+1)/2$ processors if the matrices in the second-level minimization are constructed explicitly), we can distribute the workload of the computationally intensive portion of the A2DMRG Algorithm 2, namely, steps two and three, to these processors and arrive at the following cost per processor for a single iteration:
(39) ${\rm Cost}_{\rm 1site,compute}^{\rm A2DRMG,pp}=\mathcal{O}\big(nr^{3}R(K+d)\big)+\mathcal{O}\big(dnr^{3}RK^{\prime}\big)+\mathcal{O}(d^{2}nr^{3}).$
In the specific case of quantum chemistry applications, i.e., when the energy functional $\mathcal{J}$ to be minimized is of the form (2)
$\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{\langle\mathbb{X},\mathbb{H}\mathbb{U}\rangle}{\|\mathbb{X}\|^{2}},$
with $\mathbb{H}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ denoting a second-quantized molecular Hamiltonian [19, Section 3.2], it is well-known that $R=\mathcal{O}(d^{2})$. Consequently, the cost per processor for a single iteration scales in this case as
(40) ${\rm Cost}_{\rm 1site,compute}^{\rm A2DRMG,pp}=\mathcal{O}\big(nr^{3}R(K+d)\big)+\mathcal{O}\big(dnr^{3}RK^{\prime}\big).$